OpenStudy (anonymous):

i know this sounds stupid but how do you find the value of this expression.im just starting algebra so i dont know how to do BARELY any... expression is 3x4+3

hero (hero):

3*4 + 3 = 12 + 3 = 15 It's called PEMDAS

hero (hero):

You evaluate parentheses first, then exponents, then multiplication or division, then addition or subtraction and you do it from left to right